OpenStudy (lexi724):

FAN AND MEDAL CHECK MY ANSWERS: find angle c A.70° B.75° C.80° D.85° 90-40=50 sooooo.. im lost

OpenStudy (igreen):

Picture?

OpenStudy (igreen):

The two angles at the bottom are the same, since this is an isosceles triangle. 40 + x + x = 180 40 + 2x = 180 Solve for 'x'

OpenStudy (lexi724):

subtract 2 from each side?

OpenStudy (igreen):

No, subtract 40

OpenStudy (lexi724):

to get rid of the x

OpenStudy (lexi724):

oops okay so 2x=140

OpenStudy (omarbirjas):

The answer is 60 but its not in ur answer choice....

OpenStudy (igreen):

2 is with 'x'..we don't want to do that..

OpenStudy (igreen):

Yep, you got it, now divide 2 to both sides.

OpenStudy (omarbirjas):

NEVERMIND BAD MATH

OpenStudy (lexi724):

2 would cancel and 140/2= 70

OpenStudy (lexi724):

so A

OpenStudy (omarbirjas):

Yes

OpenStudy (lexi724):

thank you!!

OpenStudy (igreen):

Yep, you got it.